OpenStudy (tylermckinney16):

You plan to rent a car to drive on your road trip. If you will be gone 5 days and plan to drive 650 miles, which of these is the better deal? Show work to explain your answer. Company A: $30 per day; unlimited mileage Company B: $15 per day; $0.10 per mile driven

OpenStudy (stamp):

While unlimited miles is a better deal if you're driving enough miles, the math reveals the true deal.

OpenStudy (triciaal):

@stamp check your B again

OpenStudy (stamp):

lol yeah moved the decimal over too many so actually 75+65>Ca and the first plan is the better deal very discerning!

OpenStudy (seratul):

Hmmm? I got the better deal is the second one.

OpenStudy (triciaal):

@TylerMckinney16 do you understand everything ? @stamp try to engage the asker so he/she will know the process

OpenStudy (seratul):

Deal 1: 30* 5 = 150 Deal 2: 15*5=75 650*.1=65 75+65=140 Deal 2 is better.

OpenStudy (triciaal):

Yes B is better by 10

OpenStudy (stamp):

yall are doing great

OpenStudy (triciaal):

@TylerMckinney16 again, any questions can you do one like this by yourself now?

OpenStudy (tylermckinney16):

Yep thanks.

OpenStudy (triciaal):

great everyone welcomes you